Find the area of the region inside the circle r=2sinθ and outside the circle r=1.
1
Expert's answer
2020-06-01T18:04:21-0400


The intersection pints of both the curves are given by :

"2 sin\theta = 1 \\
\theta = π/6 , 5π/6"


Let "r_1=1 , r_2=2sin\theta"


Area = "\frac{1}{2}\int_{π/6}^{5π/6} (r_2^2-r_1^2)d\theta\\"


"=\frac{1}{2}\int_{π/6}^{5π/6} (4sin\theta^2-1)d\theta\\
 = =\frac{1}{2}\int_{π/6}^{5π/6} (2cos2\theta+1)d\theta\\
=\frac{1}{2} (sin2\theta+\theta)|_{π/6}^{5π/6}\\
= -\frac{-\sqrt3}{2}+\frac{π}{3}"